1. (5 points) Use the limit definition to find the derivative of f(x) = 3x2 – 2+1 at x = 4. Show all steps and setup f(a+h)-f**please note the limit definition has an "a"**
Answer #1

at x=4 f(x)= 3x?- se+1 f(a) = lim f(ath) - fca) ho h f(9)= 3x 16 -4+1 = 48-4+1 = 45 f(4) = lim f (4th)-f(4) h>0 h lim £3 (4